OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? 1/2 x - 2y = 4 and 4y - x = -8

OpenStudy (a1234):

Put them into slope-intercept form: 1/2 x - 2y = 4 -2y = -1/2x + 4 y = 1/4x - 2 4y - x = -8 4y = x - 8 y = 1/4x - 2 They're exactly the same, so there are infinite solutions.

OpenStudy (anonymous):

@a1234 thank you for showing how to work it out I understand now :)

OpenStudy (a1234):

You're welcome. Glad you understand!